Please can you help me on this question: Frank earns £12 per hour. He work for 40 hours. I f he save 1/5 of his earnings each week, How many weeks will it take for him to save £500.

Found 2 solutions by rfer, balthazar612:
Answer by rfer(16322)

12*40=480
480*0.20=96
500/96=5.2 weeks

Answer by balthazar612(14)

according to statement
12*40=480
480*0.20=96
500/96=5.2 weeks answer